At Line Drive Apparel, practice jerseys are sold for $30. The disposal costs are $5 per jersey. The historical cost is $22 per jersey but the current replacement cost is $20 per jersey. What cost amount should the company use in the lower-of-cost-or-net-realizable-value (NRV) comparison if Line Drive Apparel's normal profit margin is 20% of the sale price?
A. $20
B. $22
C. $19
D. $25
Answer #1
Net-realizable-value (NRV) = 30-5 = $25
NRV less normal profit = 25-(30*20%) = $19
Market = Higher of replacement cost or NRV less normal profit
Market = Higher of $20 or $19 = $20
Inventory value = Lower of Cost or Market
Inventory value = Lower of $22 or $20 = $20
Option A $20 is correct